Neighborhood Insurance sells fire insurance policies to local homeowners. The premium is $300, the probability of a fire is 0.1%, and in the event of a fire, the insured damages (the payout on the policy) will be $290,000.

a. Make a table of the two possible payouts on each policy with the probability of each.



b. Suppose you own the entire firm, and the company issues only one policy. What are the expected value, variance and standard deviation of your profit?



c. Now suppose your company issues two policies. The risk of fire is independent across the two policies. Make a table of the three possible payouts along with their associated probabilities. (Round your "Probability" answers to 4 decimal places.)



d. What are the expected value, variance and standard deviation of your profit?



e. Compare your answers to (b) and (d). Did risk pooling increase or decrease the variance of your profit?



f. Continue to assume the company has issued two policies, but now assume you take on a partner, so that you each own one-half of the firm. Make a table of your share of the possible payouts the company may have to make on the two policies, along with their associated probabilities. (Round your "Probability" answers to 4 decimal places.)



g. What are the expected value and variance of your profit?
Answer #1
a Probability of no fire=1-0.001=0.999
NoFire Fire
Payout $0 $290,000
Probability 0.999 0.001
b Profit in case of Fire=300-290000= ($289,700)
A B C=A*B D=B-10 E=D^2 F=E*C
Probability Profit Profit*Probability Deviation fromexpected Deviation Squared Deviation Squared* Probability
Fire 0.001 ($289,700) ($289.70) ($289,710) $83,931,884,100 $83,931,884.10
No Fire 0.999 $300 $299.70 $290 $84,100 $84,015.90
Total $10.00 TOTAL $84,015,900
Expected Profit $10.00
Variance $84,015,900
Standard Deviation =Square Root of Variance
Standard Deviation $9,166 (SQRT(84015900)
c Probability of 0 Fire =0.999*0.999= 0.998001
Probability of 1 Fire =2*0.001*0.999= 0.001998 1
Probability of 2 Fire =0.001*0.001= 0.000001
Total 1
No Fire One Fire TwoFire
Payout $0 $290,000 $580,000
Probability 0.998001 0.001998 0.000001
Profit for No fire $600
Profit for One fire ($289,400) (600-290000)
Profit for two fire ($579,400)
A B C=A*B D=B-20 E=D^2 F=E*C
Probability Profit Profit*Probability Deviation fromexpected Deviation Squared Deviation Squared* Probability
No Fire 0.998001 $600 $598.80 $580 $336,400 $335,727.54
One Fire 0.001998 ($289,400) ($578.22) ($289,420) $83,763,936,400 $167,360,344.93
Two Fire 0.000001 ($579,400) ($0.58) ($579,420) $335,727,536,400 $335,727.54
TOTAL $20.00 TOTAL $168,031,800
Expected Profit $20.00
Variance $168,031,800
Standard Deviation =Square Root of Variance
Standard Deviation $12,963 (SQRT(168031800)